Reports 1-1 of 1 Clear search Modify search
fabian.arellano - 14:29 Thursday 05 September 2019 (10352) Print this report
Tripping of the HPCD of GAS stages does not trigger the tripping of the tower part.

Shoda-san and Terrence reported that the tripping triggered by the internal watchdog of the HPCD of the GAS stage does not produce a tripping of the system. The screenshots shows how the medm screens look like now.

It should be fixed soon. 

Images attached to this report
Comments to this report:
terrence.tsang - 22:04 Thursday 05 September 2019 (10362) Print this report

Turned out I misunderstood Mark's vistools library. The trippedBioWds() method in vistools returns ['BIO_GAS_MON'] instead of ['GAS'] which I was expecting and coded in the Type-B guardian. I fixed this issue and I would like to update the guardian for all Type-B suspensions on tomorrow.

INTERESTINGLY, the tripped hardware watchdog of the GAS LVDT drivers does not stop the coils from actuating. This is obvious because the integrator didn't saturated even when the hardward WD is tripped. Was the hardware watchdog of the GAS HPCD also disabled?

ayaka.shoda - 9:04 Friday 06 September 2019 (10371) Print this report

The hardware watchdog does not completely shut down the output. The actuation gain just decreases by 40 dB or so.
Therefore, the output would not saturate even when the WD is tripped if the natural offset is considerably small.
It would be possible to idnetify whether the WD works or not by looking into the output values before and after resetting the WD.

fabian.arellano - 11:17 Friday 06 September 2019 (10374) Print this report

After updating Guardian with Terrences's correction, the system was not able to go into aligned state because the actuation of the payload grew too much producing the tripping of the GAS filter coil driver. The origin of the tripping was not the payload control filters themselves but an unreliable oplev signal due to a large misalignment of the optic (klog 10373).

After fixing that problem it's possible to achive the ALIGNED state. However, the problem still to be solved is the tripping of the GAS filter coil driver when the actuation onto the optic grows too much.

fabian.arellano - 17:00 Wednesday 11 September 2019 (10464) Print this report

Today I went into the tunnel and checked the status of the binary input/output (BIO) cables for the coil drivers. The constraint was that I was not able to disconnect any cable because they were using the suspension.

According to the cable labels, some seem to be connected incorrectly. However, it's necessary to disconnect one by one in order to check the effect on each stage. The  highlights are the following:

  • The IP HPCD is labeled as #0 but the BIO cables are labeled for HPCD#1.
  • The GAS HPCD as labeled as #1 but the BIO cables are labeled for HPCD#0.
  • The two IP cable, labeled HPDC#1/DIO#2, are connected to cable without labels, cables that go to the DIO cards inside the I/O chassis.
  • One cable which is meant for HPCD#1, BIO2, DIO#2 is used for the GAS filter and IP fishing rods.
  • The other cable which si meant for HPCD#1, BIO1, DIO#2 is not connected.

Tomorrow I will aim to check again by disconnecting cables.

fabian.arellano - 10:05 Friday 13 September 2019 (10508) Print this report

Yesterday (13-09-2019) I was able to test by disconnecting cables for the coil driver binary outputs. They were incorrectly connected indeed and I fixed them.

Very likely the information will not be useful in the future but I write here how they were connected, just in case: 

Coil driver ⇒ MEDM screen indicator






Search Help